### v2

@@ 3 The Structural Complexity of Matrix-Vector Multiplication
We present a technical overview of the algorithms and the main theorems. We begin this section with a quick recap of previous work Björklund and Lingas (2001); Alves et al. (2024) who established a connection between matrix vector products and minimum spanning trees on $n$ points $\{0,1\}^{m}$, where the edge weights are given by the Hamming distance between two points (i.e., matrix $\mathbf{M}$ defines a collection of points).
Having established this connection, we prove one of our main results in Section 3.2. We prove via techniques from computational geometry that the weight of the minimum spanning tree must be bounded, if the matrix $\mathbf{M}$ has corrupted VC-dimension at most $d$. This then yields our main result Theorem 1.1.
@@ 3.1 Matrix Vector Products via Minimum Spanning Trees
Differential Compression.
Let $\mathbf{M}\in\{0,1\}^{m\times n}$ be a binary matrix, and let $\mathbf{M}_{i}$ denote the $i$’th column of $\mathbf{M}$ for $i\in[n]$. We differentially compress $\mathbf{M}$ by writing each column $\mathbf{M}_{x}$ as the sum of another column $\mathbf{M}_{y}$ and a change vector $\Delta_{x,y}$ such that $\mathbf{M}_{y}=\mathbf{M}_{x}+\Delta_{x,y}$ and $\Delta_{x,y}=\mathbf{M}_{x}-\mathbf{M}_{y}$.
If $\mathbf{M}_{x}$ and $\mathbf{M}_{y}$ are similar, then $\|\Delta_{x,y}\|_{1}$ likely to be smaller than the number of nonzero elements of $\mathbf{M}_{x}$. Therefore, it is more efficient to indirectly represent $\mathbf{M}_{x}$ with respect to $\mathbf{M}_{y}$, rather than through the matrix $\mathbf{M}$. This necessitates an algorithm that finds a suitable chain of $\Delta$’s to represent all columns of $\mathbf{M}$: for each column $\mathbf{M}_{x}$, identify a similar column $\mathbf{M}_{y}$ that characterizes the former, such that the Hamming weight of $\Delta_{x,y}$ required to represent $\mathbf{M}_{x}$ is minimized subject to $\mathbf{M}_{y}$.
Since the algorithm measures the Hamming distance for each pair of columns in $\mathbf{M}$, we can model this compression by an undirected weighted graph on $G$ on $n$ vertices, where vertex $i$ represents column $i$ of $\mathbf{M}$, and the weight of each edge $(x,y)$ corresponds to the Hamming distance between columns $x$ and $y$. We can then find a Minimum Spanning Tree (MST) of $G$, which by definition spans $G$ with the minimum sum of edge weights possible. Therefore, any MST of $G$ rooted at vertex $r$ defines a chain of $\Delta$’s that, starting at $r$, can represent all the columns of $\mathbf{M}$.
@@
For any $\Delta$-labeled spanning tree of $\mathbf{M}$ denoted by $T$ with directed edge-set $E$, define $\mathrm{weight}(T)$ as the total number of non-zero entries given by $\sum_{e\in T}\operatorname{nnz}(\Delta_{e})$.
@@
Lemma 3.3 (Gasieniec and Lingas (2003); Alves et al. (2024)).
Given a $\Delta$-labeled spanning tree $T$ of $\mathbf{M}^{\top}\in\{0,1\}^{n\times m}$ and a vector $v\in\mathbb{R}^{n}$, there is an algorithm that computes $\mathbf{M}v$ in $O(\mathrm{weight}(T)+n+m)$ time.
We describe this procedure in Algorithm 1.
As a quick recap, the algorithm computes $\mathbf{M}v$ by starting at a root node $r\in T$ of the spanning tree and performing depth-first search (DFS). The algorithm initializes an output vector and computes $(\mathbf{M}v)_{r}$ in $O(n+m)$ time. For each edge $(x,y)$ traversed in the recursion tree, the algorithm computes $(\mathbf{M}v)_{y}=(\mathbf{M}v)_{x}+\Delta_{y,x}^{\top}v$. The complexity of this iterative procedure is proportional to $\mathrm{weight}(T)=\sum_{e\in E(T)}\delta_{e}$ where for each edge $e=(x,y)$ in the $\Delta$-labeled spanning tree, $\delta_{e}$ is the number of non-zeros in $\Delta_{e}$, i.e., the Hamming distance between row $x$ and row $y$ of $\mathbf{M}$. For completeness sake, we provide a proof for Lemma 3.3 in A.7.
Since the weight of the $\Delta$-labeled spanning tree of $\mathbf{M}^{\top}$ can be different from the weight of the $\Delta$-labeled spanning tree of $\mathbf{M}$, we provide a new algorithm with an analogous theoretical guarantee in lemma 3.4.
@@
To utilize the $\Delta$-labeled spanning tree of $\mathbf{M}$, note that $\mathbf{M}v=\sum_{i=1}^{n}\mathbf{M}_{i}v_{i}$. For simplicity, consider the special case where $\mathbf{M}$ has only two columns, then $\mathbf{M}v=\mathbf{M}_{1}v_{1}+\mathbf{M}_{2}v_{2}=\mathbf{M}_{1}(v_{1}+v_{2})+\Delta_{1,2}v_{2}$. For more columns, this can be extended to require only a product with $\mathbf{M}_{1}$ and products with each $\Delta_{e}$ of the $\Delta$-labeled spanning tree. Thus the complexity is bounded by $O(m+n+\text{weight}(T))$.
We provide proofs for Lemma 3.4 in A.7.
Since we wish to minimize the time complexity, we compute a minimum spanning tree. In Alves et al. (2024) this was done naïvely, whereas in Björklund and Lingas (2001) this is done via Theorem 4.2 of Indyk and Motwani (1998).
@@
For any $\epsilon>0$, we can compute an $(1+\epsilon)$-approximate MST with respect to the Hamming distance of $n$ points in $\{0,1\}^{m}$ in $\widetilde{O}(mn^{1+1/(1+\epsilon)})$ time.
@@
Given Boolean $m\times n$ matrix $\mathbf{M}$, we can construct in $\widetilde{O}(nm)$ time a $\Delta$-labeled spanning tree whose weight is at most a $O(\log n)$-factor larger than the minimum spanning tree (MST) of $\mathbf{M}$.
Proof 3.7.
[…]
@@ 3.2 Existence of Low Weight $\Delta$-labeled Minimum Spanning Trees
The Hamming distance between any two columns in $\mathbf{M}$ is at most $m$. Since there are $O(n)$ edges in the $\Delta$-labeled MST, this gives a coarse bound of $O(mn)$ which is achieved for Hadamard matrices. One of our main technical contributions is to prove that $\Delta$-labeled MST’s with low weight exist for structure matrices. We give a bound on the MST weight parameterized by the corrupted VC-dimension of the matrix.
@@
If a matrix has corrupted VC-dimension $d$, then the weight of the corresponding MST is at most $O(mn^{1-1/d}\log n)$.
We show that for matrices of bounded VC-dimension, the minimum spanning tree $T$ has small weight. For this, we work directly with the definition of VC-dimension and use set terminology, which allows us to use relevant work from Welzl (1988); B. Chazelle (1989); Matoušek (1991); Har-Peled (2009). This previous work studied bounded VC-dimensions in the context of computational geometry. We here use their results to bound the complexity of matrix vector products.
@@
Let $A$ and $r$ be sets. We say that $A$ crosses $r$ if the symmetric difference is non-empty, i.e. $A\triangle r\neq\emptyset$. In other words, $A$ crosses $r$ if neither $A\subseteq r$ nor $A\cap r=\emptyset$ holds.
@@
Definition 3.11 (Crossing number $\kappa(T)$ of spanning tree $T$).
A crossing of a range $r\in R$ in $T$ is an edge in $T$ that is crossed by $r$ (i.e., an edge that has one endpoint in $r$ and the other endpoint in $[n]\setminus r$). The crossing number $\kappa_{r}(T)$ or a range $r\in R$ is the number of crossings of $r$ in $T$. The crossing number $\kappa(T)$ of $T$ is the maximum crossing number of a range $r\in R$ in $T$. In other words, $\kappa(T):=\max_{r\in R}\kappa_{r}(T)$.
This notion of a spanning tree corresponds to the notion of the $\Delta$-labeled spanning tree for matrix $\mathbf{M}$. The set system $(X,R)$ induced by $\mathbf{M}\in\{0,1\}^{m\times n}$ uses set $X=[n]$ (i.e., column indices) with the rows of $\mathbf{M}$ being the indicator vectors of $m$ subsets of $X$. Thus, a spanning tree $T_{X}$ for $(X,R)$ and a spanning tree $T_{M}$ for $\mathbf{M}$ both consist of edges that connect elements from $[n]$. Furthermore, the number of crossings for an edge $(x,y)\in T_{X}$ is exactly the number of indices $i$ where $(\mathbf{M}_{x})_{i}$ and $(\mathbf{M}_{y})_{i}$ differ. Thus the crossing number of a tree $T_{X}$ for $(X,R)$ is exactly the weight of the corresponding $\Delta$-labeled spanning tree for $\mathbf{M}$.
For the rest of this section, we will argue that range spaces $(X,R)$ of corrupted VC-dimension $d$ have spanning trees with small crossing number.
The literature on finding a spanning tree on $X$ with a low crossing number stems from a line of earlier works on computational geometry from the 1900s (Welzl, 1988; B. Chazelle, 1989; Matoušek, 1991; Har-Peled, 2009). They used the trees so preprocess a set of points in $\mathbb{R}^{d}$ to quickly answer queries where for any arbitrary convex polytope, the query returns a point within the polytope. Here, we transfer these techniques to the online matrix-vector multiplication problem and extend them to the notion of corrupted VC-dimension.
@@
Let $(X,R)$ be a range space VC-dimension at most $d$. For every $A\subseteq X,|A|=n\geq 1$ and every multiset $Q$ of ranges in $R$ with $|Q|=m$, there are two elements $x,y\in A$ where $x\neq y$ such that the number of ranges in $Q$ crossing $(x,y)$ is at most $O(m\cdot n^{-1/d}\log n)$.
This lemma allows us to prove the existence of a $\Delta$-labeled spanning tree for matrices with bounded VC-dimension. We later extend this to corrupted VC-dimension.
@@
Let $(X,R)$ be a range-space with $|X|=n,|R|=m$ of VC-dimension at most $d$. Then there exists a spanning tree $T$ of crossing number $\kappa(T)=O(mn^{1-1/d})$.
Proof 3.14.
Let $A_{0}=X$. Consider the following process where initially $i=0$. By lemma 3.12, there exists distinct $x_{i},y_{i}\in A_{i}$ such that the number of ranges in $R$ crossing $(x_{i},y_{i})$ is at most $O(m\cdot|A_{i}|^{-1/d}\log n)$. Letting $A_{i+1}\leftarrow A_{i}\setminus x_{i}$ and repeating, constructs a spanning tree $T$.
By lemma 3.12, the total number of ranges that cross the edges in $T$ is at most
$\sum_{i=1}^{n}O(m(n-i)^{-1/d}\log(n-i))$ $\leq O\left(m\log n\int_{1}^{n}x^{-1/d}\mathrm{d}x\right)$
$=O(mn^{1-1/d}\log n).$
which proves the claim.
We now extend Lemma 3.13 to corrupted VC-dimension. For this, we want to linearize the tree constructed in Lemma 3.13, i.e., we make sure the tree is a single line of vertices. This is to reduce the impact of corruptions, so that any single corruption does not affect the number of crossings too much.
Consider for example where $X=[n]$, $R$ is a multi-set of $m$ empty ranges (corresponding to an $m\times n$ zero matrix $\mathbf{M}$), and the spanning tree is a star with $1\in X$ at the center. The corruption matrix $\mathbf{S}$ now adds 1 to each of the empty sets in $R$ (equivalently, turning the first column of $\mathbf{M}$ to an all-1-column). This would still fit into the definition of corrupted VC-dimension since each set is corrupted by only one element. Now the number of crossings of each edge is $O(m)$ because $1$ is in each of the $m$ sets, but none of the other elements in $X$ are in any set, and there are $n-1$ edges each connecting $1$ to another element of $X$. Thus there are $O(m)$ crossings and the tree has $O(mn)$ crossings in total.
If, however, the tree had been a single line, then the corruption would affect at most 2 edges, and the number of crossings in the tree would go up by only $O(m)$.
@@
For range space $(X,R)$, $|X|=n$ with spanning tree $T$, there exists a permutation $\pi\in S_{n}$ such that the crossing number of the corresponding spanning tree $T^{\prime}$ (which is just the line $\pi(1),\pi(2),...,\pi(n)$) is at most twice that of $T$.
Proof 3.16.
The crossing number of $R_{i}$ in ${T}$ counts the number of edges in ${T}$ such that one endpoint of the edge is in $R_{i}$ and the other endpoint is in $[n]\setminus R_{i}$. Suppose the spanning tree $T$ on $X$ has a crossing number of $\kappa_{i}$ for just $R_{i}$ and let $\kappa=\sum_{i=1}^{m}\kappa_{i}$ be the crossing number of $T$ for $(X,R)$.
Perform a DFS traversal on ${T}$, and write the in-order traversal $\mathcal{I}$. We use the permutation $\pi(i)=\mathcal{I}_{i}$ for $i\in[n]$. Let $T^{\prime}$ be the corresponding linear tree.
We see then that this vertex path can (in the worst case) cross $R_{i}$ $2\kappa_{i}$ times, that is because DFS uses each edge $(x,y)$ of $T$ twice: once when visiting $y$ (going down one level in the tree) and once when it is done processing $y$ (DFS pops the current level from the stack).
Thus the total number of crossings $\kappa(T^{\prime})$ for the linear tree $T^{\prime}$ is $\kappa(T^{\prime})\leq\sum_{i=1}^{m}2\kappa_{i}\leq 2\kappa(T).$
Proof 3.17.
@@
of Theorem 3.8   Suppose that $\mathbf{M}\in\{0,1\}^{m\times n}$ has corrupted VC-dimension $d$. Let $\mathbf{L}\in\{0,1\}^{m\times n},\mathbf{S}\in\{-1,0,1\}^{m\times n}$ with $\mathbf{M}=\mathbf{L}+\mathbf{S}$ where $\mathbf{L}$ has VC-dimension $d$ and $\mathbf{S}$ has at most $O(n^{1-1/d})$ non-zero entries per row. These matrices exist by definition of corrupted VC-dimension, i.e., $\mathbf{S}$ describes the corruption.
Let $(X,R)$ be the set system induced by $\mathbf{L}$, i.e., $X=[n]$ and each row of $\mathbf{L}$ is the indicator vector of a set in $R$ with $|R|=m$. Then $(X,R)$ also has VC-dimension $d$ and by Lemma 3.13 there exists a spanning tree $T$ of $(X,R)$ with crossing number bounded by $O(mn^{1-1/d}\log n)$.
This also implies the existence of a spanning tree $T^{\prime}$ which is a single line that has crossing number $O(mn^{1-1/d}\log n)$ by Lemma 3.15.
@@
Now consider the range space $(X,R_{c})$ for the corrupted matrix $\mathbf{M}$. The spanning tree $T^{\prime}$ has the same crossing number on $(X,R_{C})$ as on $(X,R)$, except for extra crossings induced by the corruption. We consider a range $r\in R_{c}$ and count the number of new crossings it has incurred due to the corruption: an edge $(x,y)$ only has a new crossing for some range $r\in R_{c}$ if $x$ or $y$ was corrupted in $r$. Since each $x\in X$ has at most 2 incident edges in $T^{\prime}$ (since it is a line) each corruption to $r$ leads to at most $2$ new crossings. Since each range has at most $O(n^{1-1/d})$ corruptions by definition of VC-corruption, the total number of new crossing over all ranges is at most $O(mn^{1-1/d}\log n)$.
Thus, $T^{\prime}$ is a spanning tree for $(X,R_{c})$ with crossing number bounded by $O(mn^{1-1/d}\log n)$. This spanning tree corresponds to a $\Delta$-labeled spanning tree for $\mathbf{M}$, as $X=[n]$ are the column indices and the number of crossings for an edge $(x,y)$ is exactly the number of entries where $\mathbf{M}_{x}$ and $\mathbf{M}_{y}$ differ.
Since the minimum spanning tree can only have lower weight, we get an upper bound of $O(mn^{1-1/d}\log n)$ on the weight of the MST.
We can now put everything together to prove Theorem 1.1. See 1.1
Proof 3.18.
@@
Given matrix $\mathbf{M}$, preprocess it by constructing a $\Delta$-labeled spanning tree $T$ via Theorem 3.6. This takes $\widetilde{O}(mn)$ time.
When receiving a vector $v\in\mathbb{R}^{n}$, pass $v$ together with $T$ to the algorithm of Lemma 3.4. It returns the product $\mathbf{M}v$ in time $\widetilde{O}(m+n+\text{weight}(T))$.
By Theorem 3.8, if $\mathbf{M}$ has corrupted VC-dimension bounded by $d$, then the MST of $\mathbf{M}$ has weight bounded by $O(mn^{1-1/d})$. Since $T$ is a $O(\log n)$-approximation of the MST, answering a query takes $\widetilde{O}(mn^{1-1/d}+n)$ time.

### v3

@@ 3 The Structural Complexity of Matrix-Vector Multiplication
We present a technical overview of the algorithms and main theorems. We begin this section with an overview of Björklund and Lingas (2001); Alves et al. (2024) who established a connection between matrix vector products and minimum spanning trees on $n$ points $\{0,1\}^{m}$, where the edge weights are given by the Hamming distance between two points (i.e., matrix $\mathbf{M}$ defines a collection of points).
Having established this connection, we prove one of our main results in Section˜3.2. We prove via techniques from computational geometry that the weight of the minimum spanning tree is bounded if the matrix $\mathbf{M}$ has corrupted VC-dimension at most $d$. This yields our main result in Theorem˜1.3.
@@ 3.1 Matrix Vector Products via Minimum Spanning Trees
Differential Compression. Given a binary matrix $\mathbf{M}\in\{0,1\}^{m\times n}$, let $\mathbf{M}_{i}$ denote its $i$’th column. We compress $\mathbf{M}$ by writing each column $\mathbf{M}_{x}$ as the sum of another column $\mathbf{M}_{y}$ and a change vector $\Delta_{x,y}$ such that $\mathbf{M}_{y}=\mathbf{M}_{x}+\Delta_{x,y}$ and $\Delta_{x,y}=\mathbf{M}_{x}-\mathbf{M}_{y}$. If $\mathbf{M}_{x}$ and $\mathbf{M}_{y}$ are similar, then $\|\Delta_{x,y}\|_{1}$ is likely to be smaller than the number of nonzero elements of $\mathbf{M}_{x}$. So, it is more efficient to indirectly represent $\mathbf{M}_{x}$ with respect to $\mathbf{M}_{y}$, instead of through $\mathbf{M}$. This requires an algorithm that finds a chain of $\Delta$’s to represent all columns of $\mathbf{M}$: for each column $\mathbf{M}_{x}$, identify a similar column $\mathbf{M}_{y}$ such that the Hamming weight $\|\Delta_{x,y}\|_{1}$ required to represent $\mathbf{M}_{x}$ is minimized subject to $\mathbf{M}_{y}$. Since the algorithm measures the Hamming distance for each pair of columns in $\mathbf{M}$, we model this compression by an undirected weighted graph $G$ on $n$ vertices, where vertex $i$ represents $\mathbf{M}_{i}$, and the weight of each edge $(x,y)$ is $\|\Delta_{x,y}\|_{1}$. We can then find a Minimum Spanning Tree (MST) of $G$, which by definition spans $G$ with the minimum sum of edge weights possible. Therefore, any MST of $G$ rooted at vertex $r$ defines a chain of $\Delta$’s that, starting at $r$, can represent all the columns of $\mathbf{M}$.
@@
For any $\Delta$-labeled spanning tree of $\mathbf{M}$ denoted by $T$ with directed edge-set $E$, define $\mathrm{weight}(T)=\sum_{e\in T}\operatorname{nnz}(\Delta_{e})$.
@@
Lemma 3.3 (Björklund and Lingas (2001); Alves et al. (2024)).
Given a $\Delta$-labeled spanning tree $T$ of $\mathbf{M}^{\top}\in\{0,1\}^{n\times m}$ and a vector $v\in\mathbb{R}^{n}$, we can compute $\mathbf{M}v$ in $O(\mathrm{weight}(T)+n+m)$ time.
We describe this procedure in Algorithm 1. As a quick recap, the algorithm computes $\mathbf{M}v$ by starting at a root node $r\in T$ of the spanning tree and performing depth-first search (DFS). The algorithm initializes an output vector and computes $(\mathbf{M}v)_{r}$ in $O(n)$ time. For each edge $(x,y)$ traversed by DFS on the tree, the algorithm computes $(\mathbf{M}v)_{y}=(\mathbf{M}v)_{x}+\Delta_{y,x}^{\top}v$. The complexity of this iterative procedure is proportional to $\mathrm{weight}(T)=\sum_{e\in E(T)}\delta_{e}$ where for each edge $e=(x,y)$ in the $\Delta$-labeled spanning tree, $\delta_{e}$ is the number of non-zeros in $\Delta_{e}$, i.e., the Hamming distance between row $x$ and row $y$ of $\mathbf{M}$. For completeness, we prove Lemma˜3.3 in Section˜B.1.
Since the weight of the $\Delta$-labeled spanning tree of $\mathbf{M}^{\top}$ can be different from the weight of the $\Delta$-labeled spanning tree of $\mathbf{M}$, we provide a new algorithm with an analogous guarantee in Lemma˜3.4.
@@
To utilize the $\Delta$-labeled spanning tree of $\mathbf{M}$, note that $\mathbf{M}v=\sum_{i=1}^{n}\mathbf{M}_{i}v_{i}$. In the special case where $\mathbf{M}$ has only two columns, $\mathbf{M}v=\mathbf{M}_{1}v_{1}+\mathbf{M}_{2}v_{2}=\mathbf{M}_{1}(v_{1}+v_{2})+\Delta_{1,2}v_{2}$. For more columns, this can be extended to require only a product with $\mathbf{M}_{1}$ and products with each $\Delta_{e}$ in $T$. Thus the complexity is bounded by $O(m+n+\text{weight}(T))$. We prove Lemma˜3.4 in Section˜B.1.
To minimize the time complexity, we compute a MST. In Alves et al. (2024) this was done naïvely, whereas in Björklund and Lingas (2001) this is done via Theorem 4.2 of Indyk and Motwani (1998).
@@
For $\epsilon>0$, we can compute a $(1+\epsilon)$-approximate MST with respect to the Hamming distance of $n$ points in $\{0,1\}^{m}$ in time $\widetilde{O}(mn^{1+\frac{1}{1+\epsilon}})$.
@@
Given a Boolean $m\times n$ matrix $\mathbf{M}$, we can construct in $\widetilde{O}(nm)$ time a $\Delta$-labeled spanning tree whose weight is at most a $O(\log n)$-factor larger than the MST of $\mathbf{M}$.
Proof.
Run Lemma˜3.5 on the columns of $\mathbf{M}$ for $\epsilon=\log n$ to construct an $O(\log n)$-approximate MST in $\widetilde{O}(mn)$ time. Next, iterate over the edges of the MST and label each tree edge $(x,y)$ by $\Delta_{x,y}=\mathbf{M}_{y}-\mathbf{M}_{x}$. This takes $\widetilde{O}(mn)$ time as there are only n1n-1 tree edges and $\Delta_{x,y}\in\mathbb{R}^{m}$. ∎
@@ 3.2 Existence of Low Weight $\Delta$-labeled Minimum Spanning Trees
The Hamming distance between any two columns in $\mathbf{M}$ is at most $m$. Since there are $O(n)$ edges in the $\Delta$-labeled MST, this gives a coarse bound of $O(mn)$ which is achieved for Hadamard matrices. One of our main technical contributions is to prove that $\Delta$-labeled MST’s with low weight exist for structure matrices. We bound the MST’s weight parameterized by the corrupted VC-dimension of $\mathbf{M}$.
@@
If a matrix $\mathbf{M}^{\top}\in\{0,1\}^{n\times m}$ has corrupted VC-dimension $d$, then the weight the minimum spanning tree of $\mathbf{M}$ is at most $O(mn^{1-1/d}\log^{2}n)$.
We show that for matrices $\mathbf{M}$ whose transpose $\mathbf{M}^{\top}$ has bounded corrupted VC-dimension, the MST has small weight. For this, we work directly with the definition of VC-dimension and use set terminology, which allows us to use relevant work from Welzl (1988). This previous work studied bounded VC-dimensions in the context of computational geometry. We here use their results to bound the complexity of matrix vector products.
@@
Let $A$ and $r$ be sets. $r$ crosses $A$ if and only if neither $A\subseteq r$ nor $A\cap r=\emptyset$ holds. That is, there are $x,y\in A$ with $x\in r$ and $y\notin r$.
@@
Definition 3.10 (Weight of spanning tree $T$).
A crossing of a range $r\in R$ in $T$ is an edge in $T$ that is crossed by $r$ (i.e., an edge that has one endpoint in $r$ and the other endpoint in $X\setminus r$). The crossing number $\kappa_{r}(T)$ of a range $r\in R$ is the number of crossings of $r$ in $T$. The weight of tree $T$ is defined as $w(T)=\sum_{r\in R}\kappa_{r}(T)$.
This spanning tree corresponds to the $\Delta$-labeled spanning tree for matrix $\mathbf{M}$, where the range space $(X,R)$ induced by $\mathbf{M}\in\{0,1\}^{m\times n}$ has ground set $X=[n]$ (i.e., column indices) with the rows of $\mathbf{M}$ being the indicator vectors of $m$ subsets of $X$. Thus, a spanning tree $T_{X}$ for $(X,R)$ and a spanning tree $T_{M}$ for $\mathbf{M}$ both consist of edges connecting elements in $[n]$, and the number of crossings for an edge $(x,y)\in T_{X}$ equals the number of indices $i$ where $(\mathbf{M}_{x})_{i}$ and $(\mathbf{M}_{y})_{i}$ differ. Hence, $w(T_{X})$ is the weight of the corresponding $\Delta$-labeled spanning tree of $\mathbf{M}$.
We show that when the dual of $(X,R)$ has corrupted VC-dimension $d$, then $(X,R)$ has spanning trees with small weight. The literature on finding spanning trees on $X$ with few crossings stems from works in computational geometry from the 1900s (Welzl, 1988; B. Chazelle, 1989; Matoušek1991; Har-Peled, 2009) which used the trees to preprocess a set of points in $\mathbb{R}^{d}$ and answer queries where for any convex polytope, the query returns a point in the polytope. We transfer their techniques to our problem. The following lemma allows us to prove the existence of a $\Delta$-labeled spanning tree for matrices whose transpose has bounded VC-dimension. We later extend this to corrupted VC-dimension.
@@
Let $(X,R)$ be a range space with dual VC-dimension at most $d$. For every $A\subseteq X,|A|=n\geq 1$ and every multiset $Q$ of ranges in $R$ with $|Q|=m$, there exists $x\neq y\in A$ such that the number of ranges in $Q$ crossing $(x,y)$ is at most ${O}(mn^{-1/d}\log^{2}n)$.
Remark 3.12.
The original proof of Lemma 4.1 of Welzl (1988) gives a bound of $O(mn^{-1/d}\log n)$ but hides some terms that depend on the VC-dimension: the original statement of Lemma 4.1 says that the constant $c$, hidden in the $O(\cdot)$ notation, depends on $(X,R)$, but this is vacuous in our setting as $(X,R)$ is the entire problem instance. We show that $c=O(d)\leq O(\log n)$ suffices, and provide a proof in B.2 for completeness.
@@
Let $(X,R)$ be a range-space with $|X|=n,|R|=m$ with dual VC-dimension at most $d$. Then there exists a spanning tree $T$ of weight $w(T)=O(mn^{1-1/d}\log^{2}n)$.
Proof.
Let $\smash{A_{0}=X}$. Consider the following process initialized at $\smash{i=0}$. By Lemma˜3.11, there exists distinct $x_{i},y_{i}\in A_{i}$ where the number of ranges in $R$ crossing $(x_{i},y_{i})$ is $O(m|A_{i}|^{-1/d}\log^{2}n)$. Iteratively letting $A_{i+1}\leftarrow A_{i}\setminus x_{i}$ constructs a spanning tree $T$ with weight at most
$\sum_{i=1}^{n}O(m(n-i)^{-\frac{1}{d}}\log^{2}(n-i))\leq O\left(m\log^{2}n\int_{1}^{n}x^{-\frac{1}{d}}\mathrm{d}x\right)=O(mn^{1-\frac{1}{d}}\log^{2}n).\qed$
We extend Lemma˜3.13 to corrupted VC-dimensions by converting the tree to a single line of vertices, so that any single corruption minimally affects the number of crossings.
For instance, suppose $X=[n]$, $R$ is a multiset of $m$ empty ranges (corresponding to an $m\times n$ zero matrix $\mathbf{M}$), and the spanning tree is a star with $1\in X$ at the center. If the corruption adds 1 to each empty set in $R$ (equivalently, turning the first column of $\mathbf{M}$ to an all-1-column), this fits into the definition of the dual having corrupted VC-dimension. This is because each set of $(X,R)$ is corrupted by only one element, and thus in the dual set system each element is corrupted by adding it to one set.
After corruption, the number of crossings of each tree edge is $O(m)$ as $1$ is in each of the $m$ sets, but none of the other elements in $X$ are in any set. As the tree has n1n-1 edges, each connecting $1$ to another element of $X$, the tree has $O(mn)$ crossings in total. If, however, the tree was a single line, the corruption would affect at most 2 edges, and the total number of crossings would only increase by $O(m)$ rather than $O(mn)$. We convert the tree to a single line in Lemma˜3.14, proven in Section˜B.2.
@@
For a range space $(X,R)$, $|X|=n$ with spanning tree $T$, there exists a permutation $\pi\in S_{n}$ such that the weight of the corresponding spanning tree $T^{\prime}$ (which is just the line $\pi(1),\pi(2),...,\pi(n)$) is at most twice that of $T$.
Proof of Theorem˜3.7.
@@
Suppose $\mathbf{M}^{\top}=\mathbf{L}^{\top}+\mathbf{S}^{\top}\in\{0,1\}^{n\times m}$ has corrupted VC-dimension $\leq d$, where $\mathbf{L}^{\top}\in\{0,1\}^{n\times m}$ has VC-dimension $d$ and $\mathbf{S}^{\top}\in\{-1,0,1\}^{n\times m}$ has at most $O(n^{1-1/d})$ non-zero entries per column. $\mathbf{L}$ and $\mathbf{S}$ exist by definition of corrupted VC-dimension.
@@
Let $(X,R)$ be the set system induced by $\mathbf{L}$ (i.e., each row of $\mathbf{L}$ is the indicator vector of a set in $R$ with $|R|=m$), by $\mathbf{L}^{\top}$ having VC-dimension $\leq d$, $(X,R)$ has dual VC-dimension $\leq d$. By Lemma˜3.13, $(X,R)$ has a spanning tree $T$ with $w(T)\leq O(mn^{1-1/d}\log^{2}n)$. Then, by Lemma˜3.14, there exists a spanning tree $T^{\prime}$ which is a line with weight at most $O(mn^{1-1/d}\log^{2}n)$. Let $(X,R_{c})$ be the range space of $\mathbf{M}$. $T^{\prime}$ has the same weight as $(X,R_{C})$ as on $(X,R)$, excluding crossings due to the corruption. For range $r\in R_{c}$, we count the number of new crossings it incurs due to the corruption: an edge $(x,y)$ only has a new crossing for a range $r\in R_{c}$ if $x$ or $y$ was corrupted in $r$. Since each $x\in X$ has at most 2 incident edges in $T^{\prime}$ (since it is a line) each corruption to $r$ causes at most $2$ new crossings. Since the total number of corruptions is bounded by $O(mn^{1-1/d})$, we have $w(T^{\prime})\leq O(mn^{1-1/d}\log^{2}n)$ and $T^{\prime}$ is a $\Delta$-labeled spanning tree for $\mathbf{M}$, as $X=[n]$ are the column indices and the number of crossings for an edge $(x,y)$ is the number of entries where $\mathbf{M}_{x}$ and $\mathbf{M}_{y}$ differ. Since the MST has lower weight, its weight is at most $O(mn^{1-1/d}\log^{2}n)$. ∎
We can now put everything together to prove Theorem˜1.3. See 1.3
Proof.
@@
Preprocess $\mathbf{M}$ by constructing a $\Delta$-labeled spanning tree $T$ for $\mathbf{M}^{\top}$ via Theorem˜3.6 in $\widetilde{O}(mn)$ time. Given $v\in\mathbb{R}^{n}$, pass $v$ and $T$ to the algorithm of Lemma˜3.3. It returns $\mathbf{M}v$ in time $\widetilde{O}(m+n+\text{weight}(T))$. By Theorem˜3.7, the MST of $\mathbf{M}^{\top}$ has weight at most $\widetilde{O}(nm^{1-1/d})$ since $\mathbf{M}$ has corrupted VC-dimension at most $d$. Since $T$ is a $O(\log n)$-approximation of the MST, answering a query takes $\widetilde{O}(nm^{1-1/d}+m)$ time. ∎